OpenStudy (anonymous):

Use the given information to find the exact value of the expression. sin θ = 24/25, θ lies in quadrant II Find tan 2θ. A. - 526/527 B.336/625 C. -336/527 D. 336/527

OpenStudy (anonymous):

@Michele_Laino Last one help me ? /.\

OpenStudy (campbell_st):

well draw a diagram  then us the fact that \[\tan(2\theta) = \frac{2\tan(\theta)}{1 - \tan^2(\theta)}\]

OpenStudy (michele_laino):

from your data, we have that cos (theta)<0 bein theta in the II quadrant, and all points i the second quadrant have the x-coordinate <0, so: \[\cos \theta=-\sqrt{1-\frac{ 24^{2} }{ 25^{2} }}=-\frac{ 7 }{ 5 }\]

OpenStudy (campbell_st):



OpenStudy (michele_laino):

Now, we have: \[\tan \theta=\frac{ \sin \theta }{ \cos \theta }=\frac{ 24/25 }{ -7/25 }=-\frac{ 24 }{ 7 }\]

OpenStudy (anonymous):

So my answer is either a or c ?

OpenStudy (michele_laino):

Sorry, I have made an error: \[\cos \theta=-\frac{ 7 }{ 25 }\] and using the subsequent formula: \[\tan(2\theta)=\frac{ 2\tan \theta }{ 1-(\tan )^{?} }\] you get \[\frac{ 48*7 }{ 31*17 }=...\]

OpenStudy (michele_laino):

So D.

OpenStudy (anonymous):

thank you !!!!!

OpenStudy (michele_laino):

thank you!

OpenStudy (campbell_st):

if you look at the diagram a = -7 so then tan(theta) = -24/7 then using the double angle for tan \[\tan(2 \theta) = \frac{-\frac{24}{7}}{1 - (-\frac{24}{7})^2}\] just calculate it out

OpenStudy (campbell_st):

oops double the numerator

OpenStudy (michele_laino):

Sorry tI have too hurry before, the exact formula is: \[\tan(2\theta)=\frac{ 2 \tan \theta }{ 1-(\tan \theta) ^{2}}\]